OpenStudy (anonymous):

Assume the length X, in minutes, of a particular type of telephone conversation is a random variable with pdf f(x)=.2e^(-.2x), x>0. Find E[e^(.1X)+3X]. I started this problem by finding E(X)=integral x f(x) dx and with integration by parts and u substitution found the E(X) to be 5. I don't know how to proceed to find E[e^(.1X)+3X].

OpenStudy (perl):

E(x) = integral x * f(x) dx

OpenStudy (anonymous):

Thanks @perl. I actually put that in my question. I found E(X) I just don't know how to proceed to find E[e^(.1X)+3X].

OpenStudy (zarkon):

E[e^(.1X)+3X]=E[e^(.1X)]+3E[X]

OpenStudy (zarkon):

\[E[g(X)]=\int\limits_{\mathbb{R}}g(x)f(x)dx\]

OpenStudy (perl):

we can use Zarkon's suggestion, A useful formula : E[aX + b] = aE[X] + b

OpenStudy (anonymous):

So \[E[e ^{.1X}+3X] = 3(5)+e ^{.1X}\] ?

OpenStudy (perl):

$$\Large{ f(x)=.2e^{-.2x}\\ E[e^{.1X}+3X]=E[e^{.1X}]+3E[X]\\ \\ \therefore\\ = \int_{0}^{\infty}e^{.1x}(0.2e^{-.2x})dx + 3\int_{0}^{\infty}x(0.2e^{-.2x})dx \\ = \int_{0}^{\infty}0.2e^{-.1x}dx + 3(5) } $$

OpenStudy (perl):

@Zarkon why is $$ \large E[g(X)]=\int_{-\infty }^{\infty} g(x)f(x)dx $$ not $$ \large E[g(X)]=\int_{-\infty }^{\infty} g(x)f(g(x))dx $$ or are they equal

OpenStudy (perl):

it might be easier to show with a discrete p.m.f

OpenStudy (zarkon):

those two are most definitely not equal

OpenStudy (zarkon):

in general

OpenStudy (perl):

i guess im trying to replace g(x) consistently with x in the formula above. but the underlying random variable X is still the same, is that whats going on

OpenStudy (anonymous):

Thank you @zarkon and @perl for your comments. Can you tell me if in fact this should be \[=\int\limits_{0}^{\infty}e ^{.1X}(.2e ^{-.2X})dx + 3 \int\limits_{0}^{\infty}x(.2e ^{-.2x})dx\] and therefore = 2+ 3(5) = 17

OpenStudy (perl):

yes thats correct

OpenStudy (anonymous):

thank you!

OpenStudy (perl):

what if you did Y = g(X) then E [g(X)] = E[Y] = integral y * f(y)

OpenStudy (zarkon):

\[E(g(X)]=\int\limits_{\mathbb{R}}g(x)f_{X}(x)dx\] \[Y=g(X)\] \[E(Y)=\int\limits_{\mathbb{R}}yf_{Y}(y)dy\] you can prove that they are equal

OpenStudy (perl):

f_X and f_Y must be different functions. f_X is the given probability density function since you defined X = g(Y) then Y = g^-1 (X) I will work on this more later . thanks :)

OpenStudy (zarkon):

you have to be a little careful...g might not be one-to-one